OpenStudy (anonymous):

how do i do this. limit as x approaches infitnity of [e^(-7x)cos(x)]

OpenStudy (anonymous):

the instructions say evaluate

OpenStudy (anonymous):

\[\frac{\cos{x}}{e^{7x}}\]

OpenStudy (anonymous):

ohhh, small number over huge number equals negative infinity right?

OpenStudy (anonymous):

\[{x\to\infty} \implies e^{7x} \to \infty\] \(\cos\) is limited to 1 and -1, So I think Zero

OpenStudy (anonymous):

how can there be two values on top?

OpenStudy (anonymous):

I meant \(\cos\)'s Range is [-1,1]

OpenStudy (anonymous):

oh, so an infintiy under that would make it go to zero, alrite got it. thanks bro